# ATOM: A Pretrained Neural Operator for Multitask Molecular Dynamics

### v1

@@ C.5 Molecular Dynamics Simulations
We present a complete overview of the DFT parameters used to generate MD17 (Chmiela et al., 2017), RMD17 (Christensen & von Lilienfeld, 2020), MD22 (Chmiela et al., 2023), and TG80.
@@ D.1 Fully Equivariant ATOM
To achieve the full equivariance discussed in Figure 7, we employ a canonicalization network approach, which removes Euclidean gauge before learning and then reinstates it afterwards (Kaba et al., 2023). This preserves equivariance of the whole network, even with the use of non-equivariant architectures in the trunk.
@@ D.3 Value-residual Learning
In practice, we lock the first block’s $\lambda$ value to $0.5$. We report the learned $\lambda$ values in Figure 10.
(a) Aspirin
(b) Uracil
Figure 10: Learned value residuals for MD17 training over 1000 epochs.
@@ Appendix E Further Experiments
E.1 Results on tail discretization
We find the performance of both EGNO and ATOM on MD17 with tail discretization remains similar to the performance under uniform discretization discussed in table 1.
Table 8: EGNO and ATOM with final frame sampling
@@ E.1 Results on tail discretization
Gap $33.91\% +$$0.30\% +$$20.66\%$ $17.71\%$ $1.82\% +$$5.02\% +$$52.08\% +$$9.44\%$
E.2 Revised MD17 Dataset
We reach performance parity with EGNO on RMD17, shown in Table 9.
Table 9: EGNO and ATOM with final frame sampling
@@ E.2 Revised MD17 Dataset
Gap $1.47\%$ $1.27\%$ $0.88\%$ $4.39\%$ $7.21\%$ $2.78\%$ $1.23\%$ $2.00\%$
E.3 Random-split cross-validation on TG80.
For completeness, we report multitask results under compound-level random cross-validation, where compounds are randomly assigned to the train, validation, and test sets. Relative to the more challenging out-of-domain (UMAP-based) split in Table 3, EGNO is comparatively stronger; nevertheless, ATOM maintains a consistent lead across folds, with mean improvements of 24.43% on S2S and 23.93% on S2T.
@@
Gap $24.92\% +$$18.88\% +$$28.45\% +$$30.14\% +$$19.77\%$
Table 11: S2T MSE $(\times 10^{-2})$ on TG80 across five UMAP cluster assignments.
@@
Gap $27.10\% +$$15.54\% +$$27.51\% +$$35.21\% +$$14.28\%$
E.4 Multitask S2S results on TG80 under UMAP cluster cross-validation.
@@
Gap (%) $19.41\%$ $17.50\%$ $49.87\%$ $50.42\%$ $47.01\%$
E.5 Explicit Hydrogen Representation
We hypothesized that our model’s underperformance relative to EGNO in predicting malonaldehyde was due to the omission of explicit hydrogens, which limits its ability to capture electron delocalization effects involving hydroxyl hydrogens. We tested two explicit hydrogen methods:
  1. 1.
    Including all hydrogens with gradients computed for all atoms during training
  2. 2.
    Including all hydrogens but computing gradients only for heavy atoms
Contrary to conventional MD practice, neither method improved heavy-atom test loss.
Table 13: ATOM S2T MSE with implicit hydrogens (ATOM baseline) and two explicit hydrogen approaches.
Aspirin Malonaldehyde
Implicit hydrogens $6.52{\scriptstyle\pm 0.08}$ $13.51{\scriptstyle\pm 0.10}$
Explicit hydrogens 1 $7.48{\scriptstyle\pm 0.11}$ $15.15{\scriptstyle\pm 0.22}$
Explicit hydrogens 2 $6.96{\scriptstyle\pm 0.20}$ $13.52{\scriptstyle\pm 0.10}$
@@ F.1 Software and Hardware Details
All experiments were conducted using Python 3.12, NumPy 2.2.1 (Harris et al., 2020), PyTorch 2.5.1 (Paszke et al., 2019), e3nn 0.5.6 (Geiger & Smidt, 2022) and PyTorch Optimizer 3.5.0 (Kim, 2021). We use RDKit 2024.9.6 (Landrum et al., 2025) and PubChemPy 1.0.4 in the construction of TG80. All single-task training was performed on an NVIDIA RTX 5080 (16 GB) with CUDA 12.4, running on Ubuntu 24.04.
F.2 Computational Costs of Model Training
We roughly wall-clock normalised our ATOM and EGNO parameter counts, resulting in respective learnable parameter counts of 754 468 and 335 770. Despite this, ATOM still trained approximately 20% faster in single-task learning on MD17 and TG80 over three runs.
@@ F.2 Computational Costs of Model Training
EGNO Time (mins) $4.09\pm 0.15$ $3.62\pm 0.42$ $3.65\pm 0.01$ $5.01\pm 0.02$ $9.02\pm 1.22$ $5.16\pm 0.03$ $3.93\pm 0.02$ $4.35\pm 0.04$
@@
ATOM Time (mins) $5.81\pm 0.02$ $5.79\pm 0.06$ $5.79\pm 0.00$ $5.86\pm 0.01$ $5.89\pm 0.02$ $5.85\pm 0.01$ $5.81\pm 0.01$ $5.83\pm 0.02$
@@
Total FLOPS Reduction (%) $41.98\%$ $60.01\%$ $58.85\%$ $16.79\%$ $34.71\%$ $13.41\%$ $47.60\%$ $34.00\%$
In multitask training on TG80, our upsized ATOM model contained 3 557 840 parameters, compared to XXX for EGNO. Despite this, ATOM only trained between 5% and 30% slower than EGNO. This is perhaps unsurprising given the much higher FLOPS-utilization of the transformer architecture upon which ATOM is based.
@@
Total FLOPS Reduction (%) $5.72\%$ $23.29\%$ $28.49\%$ $32.99\%$ $15.71\%$
F.3 ATOM Hyperparameters
We employ the same dataset splitting and discretization parameters reported in Xu et al. (2024) for the MD17. We set the batch size to $192$, use the AdamW-AMSGrad optimizer (Loshchilov & Hutter, 2017) with an $\epsilon$ of $1\times 10^{-10}$ to avoid instability associated with the small gradients produced by zero-initialised weight matrices in early training (Jordan et al., 2025). During multitask training, we reduce the number of epochs to 250 and employ the Muon optimizer (Jordan et al., 2024; Kim, 2021). We present a complete overview of our hyperparameters in Table 16.
@@ F.3 ATOM Hyperparameters
F.4 EGNO Hyperparameters and Experimental Details
We generated the EGNO results reported in Table 1 with the same discretization parameters and hyperparameters as used in their experiments. We reduce the number of epochs from 10 00010\ 000 to 2 5002\ 500, use a batch size of 192 with the AdamW-AMSGrad optimizer (Loshchilov & Hutter, 2017), and select the best validation loss epoch for testing. In the multitask case, we further reduce the number of epochs to 250 and employ the Muon optimizer (Jordan et al., 2024; Kim, 2021). Complete hyperparameters are displayed in Table 17.
@@ G.1 Kernel Integral form of Cross-attention
We remark that the kernel fails to satisfy global Lipschitz continuity (Delattre et al., 2023), unlike FNO (Li et al., 2021), and certain generalization theorems fail as a result (Le & Dik, 2024).

### v2

@@ C.5 Molecular Dynamics Simulations
We present a complete overview of the DFT parameters used to generate MD17 (Chmiela et al., 2017), RMD17 (Christensen and von Lilienfeld, 2020), MD22 (Chmiela et al., 2023), and TG80.
@@ Appendix D Architectural Details
In this section, we discuss the finer architectural details of ATOM and the architectural adjustments made for our ablations.
@@ D.1 Fully Equivariant ATOM
To achieve the full equivariance discussed in Figure 6, we employ a canonicalization network approach, which removes Euclidean gauge before learning and then reinstates it afterwards (Kaba et al., 2023). This preserves equivariance of the whole network, even with the use of non-equivariant architectures in the trunk. This provides a controlled comparison between equivariant and non-equivariant ATOM: it enforces equivariance by construction while otherwise keeping the core architecture unchanged.
@@ D.3 Value-residual Learning
In practice, we lock the first block’s $\lambda$ value to $0.5$.
@@ Appendix E Further Experiments
We conduct further experiments on single-task and multitask learning. We consider performance under tail discretization and report results on the RMD17 dataset. For multitask learning, we report performance under random cluster assignment and S2S metrics for the experiments in Section 4.2.
E.1 Further Single-task Learning Experiments
MD17 with tail discretization We find the performance of both EGNO and ATOM on MD17 with tail discretization remains similar to the performance under uniform discretization discussed in table 1.
Table 8: EGNO and ATOM with final frame sampling. Upper part: S2S MSE. Lower part: S2T MSE.
@@ E.1 Further Single-task Learning Experiments
[…]
Revised MD17 Dataset We reach performance parity with EGNO on RMD17, shown in LABEL:tab:rmd17_results.
Table 9: EGNO and ATOM with final frame sampling. Upper part: S2S MSE. Lower part: S2T MSE.
@@
[…]
E.2 Further Multitask Learning Experiments
Random-split cross-validation on TG80. For completeness, we report multitask results under compound-level random cross-validation, where compounds are randomly assigned to the train, validation, and test sets. Relative to the more challenging out-of-domain (UMAP-based) split in Table 3, EGNO is comparatively stronger; nevertheless, ATOM maintains a consistent lead across folds, with mean improvements of 24.43% on S2S and 23.93% on S2T.
Table 10: S2S MSE $(\times 10^{-2})$ on TG80 across five random cluster assignments.
@@ E.2 Further Multitask Learning Experiments
[…]
Table 11: S2T MSE $(\times 10^{-2})$ on TG80 across five random cluster assignments.
@@
[…]
Multitask S2S results on TG80 under UMAP cluster cross-validation. The S2S side of the multitask learning results follow closely from their S2T counterparts presented in Section 4.2.
@@
[…]
Table 13: MSE $(\times 10^{-2})$ on full-dataset TG80 ATOM and single-task ATOM. S2S upper, S2T lower.
Formic Acid
Single-task ATOM 26.40
All-data ATOM 22.39
[…]
Single-task ATOM 18.10
All-data ATOM 18.72
[…]
Multitask S2S Versus Single Task We evaluate a multitask ATOM model trained on all available trajectories, pooling both ID and OOD clusters, against a single-task ATOM trained separately on each compound. The multitasking model achieves losses comparable to or lower than those of the single-task baselines, despite being trained with the same compute resources.
E.3 Monte Carlo Estimation of
Quasi-equivariance
For both the pretrained ATOM and the non-equivariant lifting variant, we estimate the quantity in Section 3.1 by Monte Carlo, drawing $N$ random timesteps $x_{n}$ and, for each, $R$ random rotations $g_{n,r}\in G$. We approximate
$\varepsilon\approx\frac{1}{N}\sum_{n=1}^{N}\left\|\frac{1}{R}\sum_{r=1}^{R}\bigl(f\bigl(\phi(g_{n,r})(x_{n})\bigr)-\rho(g_{n,r})\bigl(f(x_{n})\bigr)\bigr)\right\|_{2}.$
We report our estimates of $\varepsilon$ for various ATOM models trained on MD17 single task learning in Table 18. In all cases, ATOM shows a substantially lower equivariance defect, supporting our claim that our quasi-equivariant design achieves a kind of middle-ground in the trade-off between expressiveness and strict equivariance.
Table 14: Estimates of Quasi-equivariance $\varepsilon$ via Monte Carlo over 20 rotations and 10 timesteps with 2SD intervals.
Non-equivariant ATOM $120.23\scriptstyle\pm 153.71$ $102.50{\scriptstyle\pm 45.14}$ $26.94{\scriptstyle\pm 16.04}$ $37.97{\scriptstyle\pm 12.65}$ $37.81\scriptstyle\pm 20.08$ $57822.95\scriptstyle\pm 42872.97$ $32.29{\scriptstyle\pm 35.80}$
ATOM $\mathbf{28.34}{\scriptstyle\pm 25.10}$ $\mathbf{11.09}{\scriptstyle\pm 9.64}$ $\mathbf{9.99}{\scriptstyle\pm 9.82}$ $\mathbf{17.09}{\scriptstyle\pm 13.03}$ $\mathbf{25.46}{\scriptstyle\pm 21.47}$ $\mathbf{2744.07}{\scriptstyle\pm 3108.52}$ $\mathbf{28.72}{\scriptstyle\pm 19.20}$
@@ F.1 Software and Hardware Details
All experiments were conducted using Python 3.12, NumPy 2.2.1 (Harris et al., 2020), PyTorch 2.5.1 (Paszke et al., 2019), e3nn 0.5.6 (Geiger and Smidt, 2022) and PyTorch Optimizer 3.5.0 (Kim, 2021). We use RDKit 2024.9.6 (Landrum et al., 2025) and PubChemPy 1.0.4 to construct TG80. All single-task training was performed on an NVIDIA RTX 5080 (16 GB) with CUDA 12.4, running on Ubuntu 24.04. We use Ase 3.26.0 (Larsen et al., 2017) and MACE-Torch 0.3.14 (Kovács et al., 2025) in the experiments of Section F.3.
F.2 Training Times and Compute Requirements
Single-task training time We roughly wall-clock normalised our ATOM and EGNO parameter counts, resulting in respective learnable parameter counts of 754 468 and 335 770.
@@ F.2 Training Times and Compute Requirements
EGNO Time (mins) $4.09\pm\scriptstyle{0.15}$ $3.62\pm\scriptstyle{0.42}$ $3.65\pm\scriptstyle{0.01}$ $5.01\pm\scriptstyle{0.02}$ $9.02\pm\scriptstyle{1.22}$ $5.16\pm\scriptstyle{0.03}$ $3.93\pm\scriptstyle{0.02}$ $4.35\pm\scriptstyle{0.04}$
@@
ATOM Time (mins) $5.81\pm\scriptstyle{0.02}$ $5.79\pm\scriptstyle{0.06}$ $5.79\pm\scriptstyle{0.00}$ $5.86\pm\scriptstyle{0.01}$ $5.89\pm\scriptstyle{0.02}$ $5.85\pm\scriptstyle{0.01}$ $5.81\pm\scriptstyle{0.01}$ $5.83\pm\scriptstyle{0.02}$
@@
[…]
Multitask training time In multitask training on TG80, our upsized ATOM model contained 3 557 840 parameters, compared to 335 770 for EGNO. Despite this, ATOM only trained between 5% and 30% slower than EGNO. This is perhaps unsurprising given the much higher FLOPS-utilization of the transformer architecture upon which ATOM is based.
@@
[…]
F.3 Inference Times
We compare inference times on MD17 and MD22 across ATOM, the pretrained machine learning interaction potential MACE-OFF24 (Medium) (Kovács et al., 2025), and the classical Lennard-Jones potential (Larsen et al., 2017; Schwerdtfeger and Wales, 2024) and the molecular forcefield. We report inference times in seconds with 2SD intervals. We exclude AMBER results on Uracil as we were unable to run simulations for this molecule (Case et al., 2023).
Table 17: Seconds to produce timestep $\Delta T=3000$ of each MD17 trajectory at float32 precision.
$\Delta T=3000$ MACE-OFF $42.605\pm\scriptstyle{3.945}$ $39.656\pm\scriptstyle{0.568}$ $39.782\pm\scriptstyle{0.141}$ $39.916\pm\scriptstyle{0.155}$ $39.620\pm\scriptstyle{0.030}$ $40.271\pm\scriptstyle{2.851}$ $40.166\pm\scriptstyle{0.902}$
AMBER $37.746\pm\scriptstyle{0.536}$ $38.512\pm\scriptstyle{1.243}$ $36.806\pm\scriptstyle{0.194}$ $38.512\pm\scriptstyle{1.243}$ $37.621\pm\scriptstyle{1.412}$ $38.061\pm\scriptstyle{0.848}$ -
Lennard-Jones $2.499\pm\scriptstyle{0.266}$ $1.604\pm\scriptstyle{0.208}$ $1.529\pm\scriptstyle{0.039}$ $2.174\pm\scriptstyle{0.094}$ $1.981\pm\scriptstyle{0.011}$ $1.906\pm\scriptstyle{0.006}$ $1.804\pm\scriptstyle{0.196}$
ATOM $0.849\pm\scriptstyle{0.926}$ $0.259\pm\scriptstyle{0.097}$ $0.714\pm\scriptstyle{0.241}$ $0.467\pm\scriptstyle{0.149}$ $0.450\pm\scriptstyle{0.090}$ $0.341\pm\scriptstyle{0.075}$ $0.373\pm\scriptstyle{0.076}$
$\Delta T=10\ 000$ MACE-OFF $143.413\pm\scriptstyle{3.313}$ $142.569\pm\scriptstyle{5.584}$ $136.380\pm\scriptstyle{3.964}$ $140.288\pm\scriptstyle{9.136}$ $140.103\pm\scriptstyle{3.398}$ $140.372\pm\scriptstyle{4.754}$ $141.598\pm\scriptstyle{1.020}$
AMBER $133.900\pm\scriptstyle{3.815}$ $128.385\pm\scriptstyle{3.812}$ $121.095\pm\scriptstyle{0.968}$ $121.607\pm\scriptstyle{1.195}$ $120.319\pm\scriptstyle{1.591}$ $120.800\pm\scriptstyle{0.398}$ -
Lennard-Jones $7.797\pm\scriptstyle{0.369}$ $5.533\pm\scriptstyle{0.425}$ $5.167\pm\scriptstyle{0.221}$ $7.453\pm\scriptstyle{0.376}$ $6.721\pm\scriptstyle{0.418}$ $6.477\pm\scriptstyle{0.204}$ $5.921\pm\scriptstyle{0.522}$
ATOM $2.554\pm\scriptstyle{0.945}$ $1.002\pm\scriptstyle{0.184}$ $1.634\pm\scriptstyle{0.200}$ $098\pm\scriptstyle{0.209} 1.1$$116\pm\scriptstyle{0.083}$ $64\pm\scriptstyle{0.185} 0.990$$90\pm\scriptstyle{0.171}$
Table 18: Seconds to produce timestep $\Delta T=3000$ of each MD22 trajectory at float32 precision.
Ac-Ala3-NHME DHA Stachyose
MACE-OFF $44.737\pm\scriptstyle{4.615}$ $42.042\pm\scriptstyle{0.720}$ $47.913\pm\scriptstyle{0.546}$
Lennard-Jones $4.700\pm\scriptstyle{0.212}$ $3.790\pm\scriptstyle{0.246}$ $6.613\pm\scriptstyle{0.147}$
ATOM $1.302\pm\scriptstyle{0.766}$ $0.914\pm\scriptstyle{0.055}$ $3.075\pm\scriptstyle{0.080}$
F.4 ATOM Hyperparameters
We employ the same dataset splitting and discretization parameters reported in Xu et al. (2024) for the MD17. We set the batch size to $192$, use the AdamW-AMSGrad optimizer (Loshchilov and Hutter, 2017) with an $\epsilon$ of $1\times 10^{-10}$ to avoid instability associated with the small gradients produced by zero-initialised weight matrices in early training (Jordan et al., 2025). During multitask training, we reduce the number of epochs to 250 and employ the Muon optimizer (Jordan et al., 2024; Kim, 2021). We present a complete overview of our hyperparameters in Table 19.
@@ F.4 ATOM Hyperparameters
F.5 EGNO Hyperparameters and Experimental Details
We generated the EGNO results reported in Table 1 with the same discretization parameters and hyperparameters as used in their experiments. We reduce the number of epochs from 10 00010\ 000 to 2 5002\ 500, use a batch size of 192 with the AdamW-AMSGrad optimizer (Loshchilov and Hutter, 2017), and select the best validation loss epoch for testing. In the multitask case, we further reduce the number of epochs to 250 and employ the Muon optimizer (Jordan et al., 2024; Kim, 2021). Complete hyperparameters are displayed in Table 20.
@@ G.1 Kernel Integral form of Cross-attention
We remark that the kernel fails to satisfy global Lipschitz continuity (Delattre et al., 2023), unlike FNO (Li et al., 2021), and certain generalization theorems fail as a result (Le and Dik, 2024).